# Weyl symmetry without the traceless condition

### v3

@@ IV.2 Passive conformal transformations
may be viewed as a “rotation” of the coordinate system $R(g_{\mu\nu},\phi,\chi_{i},m)$ $\rightarrow$ $R^{\prime}(g^{\prime}_{\mu\nu},\phi^{\prime},\chi^{\prime}_{i},m^{\prime})$ (see Figure 1 of quiros_prd_2025 where, for simplicity, the vacuum case is considered). The physically meaningful quantities are the invariants:
$\mathfrak{g}_{\mu\nu}=\left(\frac{\phi}{\sqrt{6}M_{\text{pl}}}\right)^{2}\,g_{\mu\nu},\;\Psi_{i}=\left(\frac{\phi}{\sqrt{6}M_{\text{pl}}}\right)^{w_{\chi_{i}}}\chi_{i},\;\mathfrak{m}=\left(\frac{\phi}{\sqrt{6}M_{\text{pl}}}\right)^{-1}m.$ (44)
Given that these are the physical (thus measurable) quantities according to PWTs, the fields $g_{\mu\nu}$, $\phi$, $\chi_{i}$, and $m$, which are transformed according to (43) we shall call “auxiliary fields.” The gravitational state is given by ${\cal S}:(\mathfrak{g}_{\mu\nu},\Psi_{i},\mathfrak{m})$. Hence, $R$ and $R^{\prime}$ are two different “coordinate representations” of the same point $P:(\mathfrak{g}_{\mu\nu},\Psi_{i},\mathfrak{m})\in{\cal M}_{f}$; that is, of the same gravitational state $\cal S$}.
@@
is also satisfied, where $\mathfrak{D}$ is the derivative operator defined in terms of the physical metric $\mathfrak{g}_{\mu\nu}$. This equality relates the Lagrangian density of the auxiliary matter fields $\chi_{i}$ to that of the physical matter fields $\Psi_{i}$. In the Appendix A, for illustrative purposes, we show equality (45) for the Lagrangian density of the Dirac fermion and for that of perfect fluids. For other fundamental fields, e.g., gauge bosons, the demonstration is similar.
@@
where $\mathfrak{R}$ is the curvature scalar defined in terms of the Weyl invariant, thus physical, metric. Written in this (covariant) way, the total Lagrangian density coincides with that of GR. The difference between (37) and (47) according to the PWTs is that the latter Lagrangian density is written in terms of physical quantities, while the former one is written in terms of auxiliary quantities without direct physical meaning.
@@ IV.3 Active transformations
is already written covariantly in terms of physical fields $g_{\mu\nu}$, $\phi$, $\chi_{i}$, and $m$. In this case, Weyl invariants, such as those defined in (44), represent useful relationships between the physical fields in the two different gravitational states:
@@ V Gauge freedom: The “many-worlds” interpretation
The freedom to choose any arbitrary function $\phi=\phi(x)$ is called “gauge freedom”. Different gauges are fixed by different choices of the scalar $\phi=\phi(x)$. For example, the GR gauge is achieved by choosing $\hat{\phi}=\sqrt{6}\,M_{\text{pl}}=$ const. Under this choice, the total Lagrangian density (37) is written as;
@@
where $\hat{m}$ is a constant mass parameter and we took into account $\phi\rightarrow\hat{\phi}=\Omega^{-1}\phi=\sqrt{6}M_{\text{pl}}$, from which we get $\Omega=\phi/\sqrt{6}M_{\text{pl}}$. Here, we also consider the transformation $T^{(m)}_{\mu\nu}\rightarrow\hat{T}^{(m)}_{\mu\nu}=\Omega^{-2}\,T^{(m)}_{\mu\nu}$ of the matter SET; that is, $\hat{T}^{(m)}_{\mu\nu}=(\phi/\sqrt{6}M_{\text{pl}})^{-2}\,T^{(m)}_{\mu\nu}$. Since GR is a gauge of Weyl invariant theory (37), we refer to the latter theory as CGR.
Note that in an arbitrary gauge, where ${\cal L}_{m}={\cal L}_{m}(\chi_{i},\partial\chi_{i},g_{\mu\nu})$, the matter fields $\chi_{i}$ are minimally coupled to the metric $g_{\mu\nu}$. Therefore, it is the metric that defines the proper time intervals; that is, it is the physical metric in an arbitrary gauge. In the same way, the matter fields $\hat{\chi}_{i}$ are minimally coupled to the metric $\hat{g}_{\mu\nu}$, so it is the physical metric in the GR gauge. The GR gauge $\mathfrak{G}_{\text{gr}}=\left\{\sqrt{6}M_{\text{pl}},\hat{g}_{\mu\nu}(x)\right\}$, where $\hat{g}_{\mu\nu}=\hat{g}_{\mu\nu}(x)$ is any solution of the GR-EOM (51), and any other arbitrary gauge $\mathfrak{G}=\{\phi(x),g_{\mu\nu}(x)\}$, where $g_{\mu\nu}=g_{\mu\nu}(x)$ is a solution of the Einstein-type EOM (38) once we have fixed the function $\phi=\phi(x)$, represent potential descriptions of the conformal invariant gravitational interactions of matter according to CGR theory. Both solutions are related by
@@
Hence, we do not need to solve the Einstein-type EOM (38) after making a specific choice of $\phi=\phi(x)$. We only need to know some GR solution $\hat{g}_{\mu\nu}(x)$, then, if we fix the function $\phi=\phi(x)$, the corresponding conformal metric, according to (52), is necessarily a solution of (38).
@@
The usefulness of this result is illustrated, for example, in a series of papers bambi-2017-a ; bambi-2017-b ; bambi-2017-c ; bambi-2018-a ; bambi-2018-b , where conformal invariance is used in several contexts, for example, to look for the properties of some astrophysical black holes, to investigate the black hole evaporation, as well as the singularity issues. Note that in these bibliographic references, each conformal solution is assumed to be physically meaningful. However, this is true only if we follow the active approach to Weyl transformations. [^10] In contrast, if we follow the passive approach, the physical metric is the conformal invariant quantity $\mathfrak{g}_{\mu\nu}=\left(\frac{\phi}{\sqrt{6}M_{\text{pl}}}\right)^{2}\,g_{\mu\nu}=\hat{g}_{\mu\nu}.$ That is, the GR metric $\hat{g}_{\mu\nu}$ coincides with the physical metric. Therefore, only the GR solution $\hat{g}_{\mu\nu}(x)$ matters. In this case, both the GR gauge $\mathfrak{G}_{\text{gr}}$ and any arbitrary gauge $\mathfrak{G}_{k}$;
$\mathfrak{G}_{k}=\{\phi_{k}(x),g^{(k)}_{\mu\nu}(x)\},\;g^{(k)}_{\mu\nu}=\left(\frac{\sqrt{6}M_{\text{pl}}}{\phi_{k}}\right)^{2}\hat{g}_{\mu\nu},\;k=1,2,...,N,$
@@
in the conformal equivalence class $\mathfrak{C}_{\text{conf}}=\{\mathfrak{G}_{\text{gr}},\mathfrak{G}_{1},\mathfrak{G}_{2},...,\mathfrak{G}_{N}\}$ of (38), represent different gravitational states that account for different phenomenological consequences. The choice of one gauge or another can only be an experimental matter. This may be understood as a classical version of the many-worlds interpretation of quantum mechanics everett_rmp_1957 ; wheeler ; dewitt_phys_rev_1967 ; dewitt ; barvinsky_cqg_1990 ; omnes_rmp_1992 ; tegmark_1998 ; garriga_prd_2001 ; zurek_rmp_2003 ; tegmark_nature_2007 ; quiros_prd_2023 applied to the gravitational interactions of matter.
V.1 Vacuum solution: Conformal Schwarzschild space
@@
In these examples, the Schwarzschild vacuum solution $d\hat{s}^{2}$ given by (54) and its conformal singularity-free solutions; $\sqrt{6}M_{\text{pl}}/\phi(r)]^{2}\,d\hat{s}^{2}$, where $\phi(r)$ can be given by (56) or (57) (or possibly by many other $r$-functions), give suitable descriptions of vacuum CGR (55). The fact that gravitational states with a spacetime singularity are in the same equivalence class as singularity-free spacetimes is very interesting because this might shed some light on the singularity issue, as done, for example, in bambi-2017-a ; bambi-2017-b ; bambi-2017-c ; bambi-2018-a ; bambi-2018-b .
If instead of the AWTs we adopted the passive approach to WTs, the results in these bibliographic references would not have any physical significance. In fact, according to passive WTs, the physically meaningful metric $\mathfrak{g}_{\mu\nu}=[\phi(r)/\sqrt{6}M_{\text{pl}}]^{2}\,g_{\mu\nu}=\hat{g}_{\mu\nu}$, coincides with the GR vacuum metric. That is, only the Schwarzschild black hole solution (53) matters.
V.2 Cautionary note on gauge freedom
The interpretation of gauge freedom (also called gauge fixing) we have applied within the framework of CGR theory, as well as its physical meaning, has nothing in common with the usual interpretation of gauge freedom in standard gauge theories of fields. Consider, for example, the Lagrangian of a Maxwell field $A_{\mu}$ coupled to a Dirac fermion $\psi$:
where the free function $\lambda(x)$ is the gauge parameter.
In this case, gauge invariance of (58) means, in turn, that the measured quantities: the Lorentz scalar type $\bar{\psi}\psi$, vector type $\bar{\psi}\gamma^{\mu}\psi$, and tensor type $\bar{\psi}\gamma^{\mu}\gamma^{\nu}\psi$ quantities, as well as the time and spatial components of the gauge invariant field strength $F_{\mu\nu}$; the electric $E_{i}=F_{0i}$ and magnetic $B_{i}=-\epsilon_{ijk}F^{jk}/2$ fields ($\epsilon_{ijk}$ is the Levi-Civita tensor), respectively, are not affected by the gauge transformations, so $\bar{\psi}\psi=\hat{\bar{\psi}}\hat{\psi}$, $\bar{\psi}\gamma^{\mu}\psi=\hat{\bar{\psi}}\gamma^{\mu}\hat{\psi}$, …, $E_{i}=\hat{E}_{i}$ and $B_{i}=\hat{B}_{i}$. This means that these transformations do not modify the state of the Maxwell field in interaction with the fermion. This is in contrast with the interpretation of conformal transformations of the physically meaningful quantities; that is, the metric, the scalar field, and the other fields coupled to gravity, according to the active approach: $g_{\mu\nu}\rightarrow\hat{g}_{\mu\nu}=\Omega^{2}g_{\mu\nu}$, $\phi\rightarrow\hat{\phi}=\Omega^{-2}\phi$, …, where $\{\phi(x),g_{\mu\nu}(x),$...\} and $\{\hat{\phi}(x),\hat{g}_{\mu\nu}(x),$...\} represent different gravitational states, which are characterized by different curvature properties, proper time lapse, etc.
@@ VI Galactic rotation curves
In harko-prd-2023 , the possibility that the observed behavior of test particles outside galaxies, which is usually explained with the help of dark matter, could be explained, alternatively, within the framework of so-called conformal invariant Weyl geometric quadratic gravity ghilen-2019 , was investigated. The static spherically symmetric solution found in harko-epjc-2022 to the equations of motion of the latter theory was one of three cases studied in harko-prd-2023 , showing a good fit of the galactic rotation curves.
For the specific form of the Weyl vector with only a radial nonvanishing component, chosen in harko-prd-2023 ; harko-epjc-2022 , the resulting vacuum EOMs coincide exactly with (55) with the addition of a symmetry-preserving potential $V_{0}\phi^{4}/24$, for the specific choice $V_{0}=1$, once the static spherically symmetric metric (in harko-prd-2023 ; harko-epjc-2022 the BD field $\Phi=\phi^{2}$ was used):
@@
is substituted. The solution studied in harko-prd-2023 reads:
@@
Therefore, according to active CTs, the GR Schwarzschild-de Sitter metric (63), which is not suitable for explaining the galactic rotation curves, and the conformal solution (62), which can provide a good background for explaining the rotation curves, belong to the same conformal equivalence class. Both are equally plausible solutions of vacuum EOMs (55) that represent different gravitational states according to AWTs. Only experiment (in the present case, the observational data on galactic rotation curves) can choose a specific gauge among the infinity of conformally related gauges in the conformal class.
@@
If we look at the above problem from the point of view of the passive approach to CTs, as done, for instance, in hobson-epjc-2022 , there are no physical consequences of Weyl symmetry. Actually, in this bibliographic reference, it is stated that if one works in terms of quantities that can be physically measured, within the framework of conformally invariant gravity theories defined on Riemannian spacetime and having the SdS metric as a solution in the Einstein gauge, in any conformal frame, the trajectories followed by matter particles are merely the timelike geodesics of the SdS metric. Hence, there is no frame dependence of physical predictions, and it is clear that the rotation curves have no flat region. It is also stated in hobson-epjc-2022 that, since CGR is (by construction) conformally invariant, conformal transformations should not change the “observable predictions”. This makes clear that the authors consider that only conformal invariant quantities, in particular the conformal invariant metric $\mathfrak{g}_{\mu\nu}\equiv(\phi/\sqrt{6}M_{\text{pl}})^{2}g_{\mu\nu}$, have observational meaning. However, since the Weyl invariant metric $\mathfrak{g}_{\mu\nu}=(\phi/\sqrt{6}M_{\text{pl}})^{2}g_{\mu\nu}=\hat{g}_{\mu\nu}$, only the GR-SdS geometry (63) has an observational meaning. Yet, we must recall that this analysis is correct only if one assumes from the start that the passive Weyl transformations represent the only correct approach to conformal symmetry. As we have shown, if we follow this approach, the action and derived EOM of CGR theory, when written in terms of the physical quantities, coincide with the GR action and derived EOM, so Weyl symmetry is lost or is, at least, a fictitious symmetry. Recall that in this case the “auxiliary” fields that suffer the Weyl transformations: $g_{\mu\nu}$, $\phi$, $\chi_{i}$, and $m$ are removed from the covariant formulation (47).
VII Weyl invariant cosmology and the many-worlds perspective

### v4

@@ IV CGR: Physics beyond GR
may be viewed as a “rotation” of the coordinate system $R(g_{\mu\nu},\phi,\chi_{i},m)$ $\rightarrow$ $R^{\prime}(g^{\prime}_{\mu\nu},\phi^{\prime},\chi^{\prime}_{i},m^{\prime})$ (see Figure 1 of quiros_prd_2025 where, for simplicity, the vacuum case is considered). The physically meaningful quantities are the invariant metric $\mathfrak{g}_{\mu\nu}$, matter fields $\Psi_{i}$, and mass $\mathfrak{m}$:
$\mathfrak{g}_{\mu\nu}=\left(\frac{\phi}{\sqrt{6}M_{\text{pl}}}\right)^{2}\,g_{\mu\nu},\;\Psi_{i}=\left(\frac{\phi}{\sqrt{6}M_{\text{pl}}}\right)^{w_{\chi_{i}}}\chi_{i},\;\mathfrak{m}=\left(\frac{\phi}{\sqrt{6}M_{\text{pl}}}\right)^{-1}m,$ (41)
respectively. We have that $\mathfrak{g}_{\mu\nu}$, $\Psi_{i}$, and $\mathfrak{m}$ are the physical (thus measurable) quantities according to PWTs, while the fields $g_{\mu\nu}$, $\phi$, $\chi_{i}$, and $m$, which are transformed according to (40) are just “auxiliary fields.” The global gravitational state is given by ${\cal S}:(\mathfrak{g}_{\mu\nu},\Psi_{i},\mathfrak{m})$. Hence, $R$ and $R^{\prime}$ are two different “coordinate representations” of the same point $P:(\mathfrak{g}_{\mu\nu},\Psi_{i},\mathfrak{m})\in{\cal M}_{f}$; that is, of the same GGS $\cal S$}.
@@
is also satisfied, where $\mathfrak{D}$ is the derivative operator defined in terms of the physical metric $\mathfrak{g}_{\mu\nu}$. This equality relates the Lagrangian density of the auxiliary matter fields $\chi_{i}$ to that of the physical matter fields $\Psi_{i}$. In quiros-arxiv-2025 , equality (42) is shown for the Lagrangian density of the Dirac fermion and for that of perfect fluids. [^12] For perfect fluids, the demonstration of (42) is quite simple. The Lagrangian density for a perfect fluid with energy density $\rho$ can be written as ${\cal L}_{\text{fluid}}[\rho,g]=-\sqrt{-g}\,\rho$. If we define the Weyl invariant energy density $\mathfrak{r}:=(M_{\text{pl}}/\sqrt{6}\phi)^{4}\rho$, and take into account the definition of the Weyl invariant metric tensor (41), then, since $\sqrt{-\mathfrak{g}}=(\sqrt{6}\phi/M_{\text{pl}})^{4}\sqrt{-g}$, we get that $\sqrt{-\mathfrak{g}}\,\mathfrak{r}=\sqrt{g}\,\rho$; that is, ${\cal L}_{\text{fluid}}[\rho,g]={\cal L}_{\text{fluid}}[\mathfrak{r},\mathfrak{g}]$, which is (42) for the case of perfect fluids. For other fundamental fields, e. g., gauge bosons, the demonstration is similar.
@@
where $\mathfrak{R}$ is the curvature scalar defined in terms of the Weyl-invariant, thus physical, metric. Written in this (covariant) way, the total Lagrangian density coincides with that of GR. The difference between (24) and (44) according to the PWTs is that the latter Lagrangian density is written in terms of physical quantities, while the former one is written in terms of auxiliary fields without direct physical meaning.
@@
is already written covariantly in terms of physical fields $g_{\mu\nu}$, $\phi$, $\chi_{i}$, and $m$. In this case, Weyl invariants, such as those defined in (41), represent useful relationships between the physical fields in the two different global gravitational states:
@@ V Gauge freedom: The “many-worlds” interpretation
The freedom to choose any arbitrary function $\phi=\phi(x)$ is called “gauge freedom”. Different gauges are fixed by different choices of the scalar $\phi=\phi(x)$. For example, the GR gauge is achieved by choosing $\hat{\phi}=\sqrt{6}\,M_{\text{pl}}=$ const. [^13] We underline that once we make an specific gauge choice $\phi=\phi(x)$, the explicit Weyl invariance is lost. This is similar to the loss of explicit coordinate invariance after a specific choice of coordinates is made within the GR framework. Under this choice, the total Lagrangian density (24) is written as;
@@
where $\hat{m}$ is a constant mass parameter and we took into account $\phi\rightarrow\hat{\phi}=\Omega^{-1}\phi=\sqrt{6}M_{\text{pl}}$, from which we get $\Omega=\phi/\sqrt{6}M_{\text{pl}}$. Here, we also consider the transformation $T^{(m)}_{\mu\nu}\rightarrow\hat{T}^{(m)}_{\mu\nu}=\Omega^{-2}\,T^{(m)}_{\mu\nu}$ of the matter SET; that is, $\hat{T}^{(m)}_{\mu\nu}=(\phi/\sqrt{6}M_{\text{pl}})^{-2}\,T^{(m)}_{\mu\nu}$. Since GR is a gauge of Weyl-invariant theory (24), we refer to the latter theory as CGR.
Note that in an arbitrary gauge, where ${\cal L}_{m}={\cal L}_{m}(\chi_{i},\partial\chi_{i},g_{\mu\nu})$, the matter fields $\chi_{i}$ are minimally coupled to the metric $g_{\mu\nu}$. Therefore, it is the metric that defines the proper time intervals; that is, it is the physical metric in that gauge. In the same way, the matter fields $\hat{\chi}_{i}$ are minimally coupled to the metric $\hat{g}_{\mu\nu}$, so it is the physical metric in the GR gauge. Let us take, for example, the GR gauge $\mathfrak{G}_{\text{gr}}=\left\{\sqrt{6}M_{\text{pl}},\hat{g}_{\mu\nu}(x),\hat{\chi}_{i},\hat{m}\right\}$, where $\hat{g}_{\mu\nu}=\hat{g}_{\mu\nu}(x)$ is any solution of the GR-EOM (48) and $\hat{m}$= const. is the GR constant mass parameter, and any other arbitrary gauge $\mathfrak{G}=\{\phi(x),g_{\mu\nu}(x),\chi_{i},m\}$, where $g_{\mu\nu}=g_{\mu\nu}(x)$ is a solution of the Einstein-type EOM (33) once we have fixed the function $\phi=\phi(x)$ (it is assumed that the matter fields are given). These gauges represent potential descriptions of the conformal invariant gravitational interactions of matter according to CGR theory, where the respective solutions are related by
@@
Hence, we do not need to solve the Einstein-type EOM (33) once we make a specific choice for $\phi=\phi(x)$. We only need to know some GR solution $\hat{g}_{\mu\nu}(x)$, then, if we fix the function $\phi=\phi(x)$, the corresponding conformal metric, according to (49), is necessarily a solution of (33).
@@
The usefulness of this result is illustrated, for example, in a series of papers bambi-2017-a ; bambi-2017-b ; bambi-2017-c ; bambi-2018-a ; bambi-2018-b , where Weyl invariance is used in several contexts, for example, to look for the properties of some astrophysical black holes, to investigate the black hole evaporation, as well as the singularity issues. Note that in these bibliographic references, each conformal solution is assumed to be physically meaningful. However, this is true only if we follow the active approach to Weyl transformations. [^14] In contrast, if we follow the passive approach, the physical metric is the conformal invariant quantity $\mathfrak{g}_{\mu\nu}=\left(\frac{\phi}{\sqrt{6}M_{\text{pl}}}\right)^{2}\,g_{\mu\nu}=\hat{g}_{\mu\nu}.$ That is, the GR metric $\hat{g}_{\mu\nu}$ coincides with the physical metric. Therefore, only the GR solution $\hat{g}_{\mu\nu}(x)$ matters. In this case, the GR gauge $\mathfrak{G}_{\text{gr}}$ and any arbitrary gauge $\mathfrak{G}_{k}$;
$\mathfrak{G}_{k}=\{\phi_{k}(x),g^{(k)}_{\mu\nu}(x),...\},\;g^{(k)}_{\mu\nu}=\left(\frac{\sqrt{6}M_{\text{pl}}}{\phi_{k}}\right)^{2}\hat{g}_{\mu\nu},\;k=1,2,...,N,$ (50)
where $N\rightarrow\infty$ and the ellipsis within curled brackets stands for other fields coupled to gravity, belong in the conformal equivalence class
$\mathfrak{C}_{\text{conf}}=\{\mathfrak{G}_{\text{gr}},\mathfrak{G}_{1},\mathfrak{G}_{2},...,\mathfrak{G}_{N}\},$ (51)
of CGR theory given by the Lagrangian density (24) and the derived EOM (33). As a result, these gauges correspond to different global gravitational states that describe distinct phenomenological consequences. The choice between one gauge and another is ultimately an experimental matter.
@@
The arising picture may be understood as a classical version of the many-worlds interpretation of quantum mechanics everett_rmp_1957 ; wheeler ; dewitt_phys_rev_1967 ; dewitt ; barvinsky_cqg_1990 ; omnes_rmp_1992 ; tegmark_1998 ; garriga_prd_2001 ; zurek_rmp_2003 ; tegmark_nature_2007 ; quiros_prd_2023 applied to the gravitational interactions of matter. Besides, in the present case, the many worlds picture arises not in spacetime but in the configuration space.
V.1 Cautionary note on gauge freedom
The interpretation of gauge freedom that we have applied within the framework of CGR theory, as well as its physical meaning, has nothing in common with the usual interpretation of gauge freedom in standard gauge theories of fields. Consider, for example, the Lagrangian of a Maxwell field $A_{\mu}$ coupled to a Dirac fermion $\psi$:
where the free function $\lambda(x)$ is the gauge parameter.
In this case, gauge invariance of (52) means that the derived EOM, as well as the measured quantities; that is, the physically meaningful quantities: the Lorentz scalar type $\bar{\psi}\psi$, vector type $\bar{\psi}\gamma^{\mu}\psi$, tensor type $\bar{\psi}\gamma^{\mu}\gamma^{\nu}\psi$ quantities, and the time and spatial components of the gauge invariant field strength $F_{\mu\nu}$; the electric $E_{i}=F_{0i}$ and magnetic $B_{i}=-\epsilon_{ijk}F^{jk}/2$ fields ($\epsilon_{ijk}$ is the Levi-Civita tensor), respectively, are not affected by the gauge transformations; that is, $\bar{\psi}\psi=\hat{\bar{\psi}}\hat{\psi}$, $\bar{\psi}\gamma^{\mu}\psi=\hat{\bar{\psi}}\gamma^{\mu}\hat{\psi}$, …, $E_{i}=\hat{E}_{i}$ and $B_{i}=\hat{B}_{i}$. This means that these transformations do not modify the state of the Maxwell field in interaction with the fermion. Note that, in the present case, the fields $A_{\mu}$, $\psi$, and $\bar{\psi}$, by themselves, do not represent measured quantities, so these are physically meaningless. Even if we try to associate different electromagnetic (EM) states with different sets of these fields: $\mathfrak{E}:(\psi,\bar{\psi},A_{\mu})$, $\hat{\mathfrak{E}}:(\hat{\psi},\hat{\bar{\psi}},\hat{A}_{\mu})$, etc., these states share the same measured quantities so that they are physically undistinguishable. To all practical purposes, $\mathfrak{E}$ and $\hat{\mathfrak{E}}$ represent the same physical EM state. This is in contrast with the active interpretation of Weyl transformations of the physically meaningful quantities; that is, the metric, the scalar field, and the other fields coupled to gravity: $g_{\mu\nu}\rightarrow\hat{g}_{\mu\nu}=\Omega^{2}g_{\mu\nu}$, $\phi\rightarrow\hat{\phi}=\Omega^{-2}\phi$, …, where $\{\phi(x),g_{\mu\nu}(x),$...\} and $\{\hat{\phi}(x),\hat{g}_{\mu\nu}(x),$...\} represent different global gravitational states, which are characterized by different measured quantities; that is, different curvature properties, proper time lapse, etc.
V.2 Vacuum solution: Conformal Schwarzschild space
@@
In these examples, the Schwarzschild vacuum solution $d\hat{s}^{2}$ given by (54) and its conformal singularity-free solutions; $\sqrt{6}M_{\text{pl}}/\phi(r)]^{2}\,d\hat{s}^{2}$, where $\phi(r)$ can be given by (56) or (57) (or possibly by many other $r$-functions), give suitable descriptions of vacuum CGR (55). The fact that global gravitational states with a spacetime singularity are in the same equivalence class as singularity-free spacetimes is very interesting because this might shed some light on the singularity issue, as done, for example, in bambi-2017-a ; bambi-2017-b ; bambi-2017-c ; bambi-2018-a ; bambi-2018-b . [^17] If instead of the AWTs we adopted the passive approach to WTs, the results in bambi-2017-a ; bambi-2017-b ; bambi-2017-c ; bambi-2018-a ; bambi-2018-b would not have any physical significance. In fact, according to passive WTs, the physically meaningful metric $\mathfrak{g}_{\mu\nu}=[\phi(r)/\sqrt{6}M_{\text{pl}}]^{2}\,g_{\mu\nu}=\hat{g}_{\mu\nu}$, coincides with the GR vacuum metric. That is, only the Schwarzschild black hole solution (53) matters. Yet, it is ultimately the experiment that decides which gauge is the one that better describes the gravitational phenomena. If it were, for example, the GR gauge, the one selected by experiment, then the spacetime singularity would still be there. Nevertheless, as we will show in Section IX, consideration of quantum-mechanical effects in the present framework may remove the classical spacetime singularity.
@@ VI Galactic rotation curves
The possibility that the observed behavior of test particles outside galaxies, which is usually explained with the help of dark matter, could be explained, alternatively, within the framework of so-called conformal invariant Weyl geometric quadratic gravity ghilen-2019 , was investigated in harko-prd-2023 . The static spherically symmetric solution found in harko-epjc-2022 of the EOM of the latter theory was one of three cases studied in harko-prd-2023 , showing a good fit to the galactic rotation curves.
For the specific form of the Weyl vector with only a radial nonvanishing component, chosen in harko-prd-2023 ; harko-epjc-2022 , the resulting vacuum EOMs coincide exactly with (55) with the addition of a Weyl symmetry-preserving potential $\phi^{4}/24$ (in harko-prd-2023 ; harko-epjc-2022 the BD field $\Phi=\phi^{2}$ was used). Then, there is no doubt that the solution found in these publications is a vacuum solution of Weyl-invariant CGR theory (24) with a cosmological constant. As we shall see, the solution involves a specific function $\phi=\phi(r)$, so that it is in fact a gauge of vacuum CGR theory (55) with a cosmological constant. Let us discuss the mentioned solution that provides a good explanation of the galaxies’ rotation curves.
In harko-prd-2023 ; harko-epjc-2022 the static spherically symmetric metric
@@
was assumed. The solution studied in harko-prd-2023 reads:
@@
This is an expected result, since (59) is a vacuum solution of CGR with a cosmological constant. This means that GR $\Lambda$ is a gauge of this theory when we set $\phi(r)=\sqrt{6}M_{\text{pl}}$. Hence, any solution (61) that is conformal to the Schwarzschild-de Sitter vacuum solution (62) is also a solution of (55) plus a cosmological constant term.
@@
According to active CTs, the GR Schwarzschild-de Sitter metric (62), which is not suitable for explaining the galactic rotation curves, and the conformal solution (61), which can provide a good background for explaining the rotation curves, belong to the same conformal equivalence class $\mathfrak{C}_{\text{conf}}$ (51). Both are equally plausible solutions of vacuum EOM (55) (plus a $\Lambda$ term) that represent different global gravitational states according to AWTs. Only experiment (in the present case, the observational data on galactic rotation curves) can choose a specific gauge among the infinity of conformally related gauges in the conformal class $\mathfrak{C}_{\text{conf}}$ of vacuum CGR theory (plus a $\Lambda$ term).
@@
If we look at the above problem from the point of view of the passive approach to CTs, as done, for instance, in hobson-epjc-2022 , there are no physical consequences of Weyl symmetry. Actually, in this bibliographic reference, it is stated that if one works in terms of quantities that can be physically measured, within the framework of conformally invariant gravity theories defined on Riemannian spacetime and having the SdS metric as a solution in the Einstein gauge, in any conformal frame, the trajectories followed by matter particles are merely the timelike geodesics of the SdS metric. Hence, there is no frame dependence of physical predictions, and it is clear that the rotation curves have no flat region. It is also stated in hobson-epjc-2022 that, since CGR is (by construction) conformally invariant, conformal transformations should not change the “observable predictions”. This makes clear that the authors consider that only conformal invariant quantities, in particular the conformal invariant metric $\mathfrak{g}_{\mu\nu}\equiv(\phi/\sqrt{6}M_{\text{pl}})^{2}g_{\mu\nu}$, have observational meaning. However, since the Weyl-invariant metric $\mathfrak{g}_{\mu\nu}=(\phi/\sqrt{6}M_{\text{pl}})^{2}g_{\mu\nu}=\hat{g}_{\mu\nu}$, only the GR-SdS geometry (62) has an observational meaning. However, we must recall that this analysis is correct only if one assumes from the start that the passive Weyl transformations represent the only correct approach to conformal symmetry. As we have shown, if we follow this approach, the action and derived EOM of CGR theory, when written in terms of the physical quantities, coincide with the GR action and derived EOM, so Weyl symmetry is lost or is, at least, a fictitious symmetry. Recall that in this case the “auxiliary” fields that suffer the Weyl transformations: $g_{\mu\nu}$, $\phi$, $\chi_{i}$, and $m$ are removed from the covariant formulation (44).
VII Weyl-invariant cosmology and the many-worlds perspective